# Video LLMs for Temporal Reasoning in Long Videos

### v4

@@ Abstract
This paper introduces TemporalVLM, a video large language model (video LLM) capable of effective temporal reasoning and fine-grained understanding in long videos. At the core, our approach includes a visual encoder for mapping a long-term input video into features which are time-aware and contain both local and global cues. In particular, it first divides the input video into short-term clips, which are jointly encoded with their timestamps and fused across overlapping temporal windows into time-sensitive local features. Next, the local features are passed through a bidirectional long short-term memory (BiLSTM) module for global feature aggregation. The extracted time-aware and multi-level features are important for accurate temporal reasoning and fine-grained understanding in long videos. Moreover, to facilitate the evaluation of TemporalVLM, we present a large-scale long video dataset of industry assembly processes, namely IndustryASM, which consists of videos recorded on factory floors with actions and timestamps annotated by industrial engineers for time and motion studies and temporal action segmentation evaluation. Finally, extensive experiments on datasets of long videos, including TimeIT and IndustryASM, show that TemporalVLM achieves superior performance than previous methods across temporal reasoning and fine-grained understanding tasks, namely dense video captioning, temporal video grounding, video highlight detection, and temporal action segmentation. To the best of our knowledge, our work is the first to incorporate LSTMs into video LLMs.
Figure 1: Previous video large language models (a-c) usually are not time-sensitive (a, b), consider an input video as a single clip (a, c), and apply pooling operation (a, b) or query aggregation (c) for aggregating global semantic information. In contrast, our model (d) includes a time-aware clip encoder, which extracts time-aware fine-grained cues from short-term clips sampled from a long-term input video, and a bidirectional long short-term memory module, which captures long-range temporal dependencies across multiple clips. Our extracted time-aware and multi-level features are crucial for temporal reasoning in long videos.
@@ 1 Introduction
Video temporal reasoning represents the process of reasoning about time and its passage in videos, with a focus on how events or actions happen and relate to each other in terms of time. Several video understanding applications require the ability of temporal reasoning. Examples include dense video captioning [46, 60, 53], which attempts to detect the timestamps and generate detailed descriptions of all the events or actions in an input video, and temporal action segmentation [23, 54, 2, 26], which aims to associate each frame of a long video capturing a complex activity with one of the action/sub-activity classes. Considerable efforts [53, 27, 30, 26] have been invested in developing separate models for solving individual tasks. These models usually have different architectures. Therefore, it is favorable to design a unified model for handling various tasks.
@@
The past few years have witnessed the impressive instruction following ability and remarkable comprehension and generation capability of large language models (LLMs) [1, 4, 41, 31, 42, 43], which have emerged as a universal agent for performing various tasks. Video LLMs [22, 24, 28, 29, 57, 16, 39] which incorporate video encoders with LLMs for video understanding tasks have been introduced. These methods often represent a video by a fixed number of visual tokens, leading to reduced performance with long videos, and encode video frames and their timestamps separately and hence struggle to perform temporal reasoning tasks. Recently, a few works [33, 13, 14, 36] have developed video LLMs with temporal reasoning abilities, e.g., TimeChat [36] proposes to vary the number of final video tokens based on the input video length and jointly encode video frames and their timestamps. The above methods usually treat the entire video as a single clip [33, 13, 14, 36] and aggregate video tokens via pooling operation [14] and query aggregation [36], struggling to capture fine-grained information, especially for long videos.
@@
We propose TemporalVLM, a video LLM capable of effective temporal reasoning and fine-grained understanding in long videos. Fig. 1 shows the differences in model architectures of TemporalVLM and previous works. Our model includes two key architectural contributions for temporal reasoning in long videos. Firstly, we divide a long-term input video into multiple short-term clips and introduce a time-sensitive clip encoder for extracting fused time-aware local features from each clip. Secondly, we adopt a bidirectional long short-term memory (BiLSTM) module which takes all the local features as inputs and computes global features from multiple clips. As a consequence, our features not only are time-sensitive but also contain both local fine-grained and global semantic information, which are crucial for temporal reasoning in long videos. Moreover, to further evaluate our model, we present IndustryASM, a large-scale long video dataset of industry assembly processes for temporal action segmentation benchmarking and time and motion studies. Our IndustryASM dataset comprises of $4803$ videos with an average video duration of $119$ seconds. It covers in total $52$ diverse industry assembly tasks and includes timestamp and action labels. We convert the labels into chat samples with manually written instructions. To the best of our knowledge, it is the first large-scale video dataset recorded on factory floors and labeled by industrial engineers for temporal action segmentation evaluation and time and motion analyses. Lastly, extensive experiments on datasets of long videos, i.e., TimeIT [36] and IndustryASM, demonstrate that our TemporalVLM approach achieves superior results over previous methods across temporal reasoning and fine-grained understanding tasks, namely dense video captioning, temporal video grounding, video highlight detection, and temporal action segmentation.
@@
  • We develop TemporalVLM, a video LLM for temporal reasoning and fine-grained understanding in long videos. Given a long-term input video, we first sample short-term clips and encode each clip along with its timestamps into fused time-aware local features. The local features are then fed to a BiLSTM module to extract global features. Therefore, our features not only are time-sensitive but also capture effective local and global information.
  • We present IndustryASM, a large-scale long video dataset of manufacturing assembly procedures. Our videos are recorded on factory floors with actions and timestamps labeled by industrial engineers for time and motion analyses and temporal action segmentation evaluation. We will make our dataset available publicly.
  • Our evaluations show that TemporalVLM outperforms prior works across temporal reasoning and fine-grained understanding tasks. To our best knowledge, this is the first work to integrate LSTMs into video LLMs.
@@ 2 Related Work
Large Language Models. Over the past few years, LLMs [1, 4, 41, 31, 42, 43] have revolutionized the field of natural language processing. Initial methods employ encoder-decoder architectures [17, 35], whereas decoder-only models, e.g., GPT [34], demonstrate impressive performance and generalization in recent years. In addition, models with billions of parameters, e.g., GPT-4 [1], have been developed to push the boundary further, while instruction tuning techniques [32, 47, 48] have been widely used to enable models to generate contextually relevant responses. Here, we focus on video LLMs, which take videos as additional inputs and explore the use of LLMs for video understanding tasks.
@@
Video Large Language Models. Motivated by the impressive reasoning and generalization abilities of LLMs, video LLMs [22, 24, 28, 29, 57, 16, 39] have been developed for video understanding tasks. They typically include three key components: i) a pre-trained visual encoder (e.g., ViT [38]) to extract visual features, ii) a projection layer to map visual features into the text latent space of LLMs, and iii) a pre-trained LLM (e.g., LLaMA [42]) for generating responses. They mostly differ in their visual encoder, i.e., how visual features are extracted and aggregated from an input video. For example, VideoChat [22] extracts frame features via a visual transformer [38] and employs a query transformer (Q-Former) [21] to aggregate frame features into video features, while a video Q-Former is further included for temporal modeling in Video-LLaMA [57]. The above methods usually map the video into a fixed number of visual tokens, yielding degrading performance with long videos, while encoding video frames and their timestamps separately and hence struggling with temporal reasoning tasks. To address these drawbacks, methods with temporal reasoning capabilities have been introduced, e.g., TimeChat [36] presents a sliding video Q-Former to handle various video lengths and a time-aware frame encoder to jointly encode video frames and their timestamps. These methods often consider the video as a single clip [28, 29, 22, 57] and aggregate video tokens via pooling operation [28, 29] and query aggregation [22, 57], overlooking fine-grained details. In this work, we divide the video into multiple clips and propose a time-aware clip encoder for capturing local features from each clip. Also, we introduce a BiLSTM module for aggregating global features over multiple clips. To our best knowledge, our work is the first to exploit LSTMs for video LLMs.
@@
Video Temporal Reasoning. Temporal reasoning plays an important role in many video understanding tasks, e.g., dense video captioning [46, 60, 53], temporal video grounding [49, 27], video highlight detection [20, 30], and temporal action segmentation [23, 54, 2, 26]. Previous works [53, 27, 30, 26] often focus on designing separate models for tackling individual tasks, while LLM-based models [33, 13, 14, 36] capable of handling multiple tasks have emerged recently. Our TemporalVLM model belongs to the second group. Also, we introduce IndustryASM, a large-scale long video dataset of industry assembly processes for temporal action segmentation and time and motion analyses.
@@
Long Video Understanding. There exist a few challenges in long video understanding such as complex spatial-temporal relationships and redundant information. Several methods have been developed for long video understanding, including efficient architectures [6, 18], temporal pooling/aggregation [37, 51], and dynamic clip selection [9, 10]. For vision-language understanding tasks, methods based on temporal alignment [11, 3] have been introduced, while memory techniques [52, 58] are often used in dense video prediction tasks. Recently, Weng et al. [50] propose LongVLM for long video understanding. It divides the video into clips and employs a merging module for extracting local features and pooling operation for computing global features from multiple clips. It does not utilize timestamps, which are crucial for temporal reasoning. In contrast, our TemporalVLM model explicitly utilizes timestamps via a time-aware clip encoder, yielding time-aware features. Moreover, we employ a learnable BiLSTM module for aggregating global features from multiple clips.
3 TemporalVLM: A Video LLM for Temporal Reasoning in Long Videos
Figure 2: Given a long-term input video with timestamps, we divide it into short-term clips with timestamps. The time-aware clip encoder is employed on each clip along with its timestamps for extracting fused time-aware local features, which capture local fine-grained cues within each clip. Next, the bidirectional long short-term memory module is adopted for computing global features, which capture global temporal relationships across multiple clips. Lastly, the large language model takes video and text tokens as inputs and generates responses.
Below we describe our first contribution, TemporalVLM, a video LLM for effective temporal reasoning in long videos. Two novel components are proposed in TemporalVLM: i) a time-aware clip encoder for extracting time-aware local features from a short-term clip, and ii) a BiLSTM module for computing global features from the local features. We provide an overview of our TemporalVLM model in Fig. 2.
3.1 Time-Aware Clip Encoder
@@
Video LLMs often consider a long-term input video as a single clip and uniformly sample a fixed number of frames $N^{v}_{f}$ from the video, e.g., $N^{v}_{f}=96$ frames in TimeChat [36] and $N^{v}_{f}=100$ frames in VTimeLLM [13]. Moreover, prior works usually use query aggregation, e.g., TimeChat [36], and pooling operation, e.g., Video-ChatGPT [29], to aggregate video tokens. Thus, they often perform well with short-term videos but miss out important fine-grained information for long-term videos, leading to degrading performance.
@@
We propose a time-aware clip encoder (i.e., blue block in Fig. 2) for extracting local fine-grained cues within a clip. Given a long-term input video, we first decompose it into $C$ short-term clips and uniformly sample a fixed number of frames $N^{c}_{f}$ from each clip. Here, we set $C=6$ clips and $N^{c}_{f}=96$ frames per clip. The sampled frames for each clip along with their corresponding timestamps are then passed to our time-aware clip encoder for simultaneously encoding the frame contents and their timestamps, yielding fused time-aware local features. The time-aware frame encoder uses a pre-trained image encoder [40] to obtain frame features, which are jointly encoded with timestamps via an image Q-Former [5], yielding time-aware frame features $\mathbf{f}_{t}$. These frame features $\mathbf{f}_{t}$ are then fed to a video Q-Former  [36] across overlapping windows $\mathbf{W}_{i}$ of size $q$ = 32 and overlap $o$ = 16 for temporal modeling. The window-wise outputs $\mathbf{V}_{i}$ are then concatenated into $\mathbf{S}$ :
[…]
$\mathbf{S}$ $=[\mathbf{V}_{1},\mathbf{V}_{2},\dots,\mathbf{V}_{n}],$ (2)
@@
with $n$ denoting the number of overlapping windows within a short-term clip. We propose a fusion module to align diverse temporal cues from redundant boundary tokens in $\mathbf{S}$. Specifically, we apply multiheaded self-attention on $\mathbf{S}$, yielding $\mathbf{C}$ which fuses local contexts across multiple windows into a single context-aware embedding as:
@@
$\mathbf{C}$ $=[\mathbf{C}^{(1)},\mathbf{C}^{(2)},\dots,\mathbf{C}^{(H)}]\mathbf{W}^{O}$ (4)
We first project $\mathbf{S}$ into $\mathbf{Q}$, $\mathbf{K}$, $\mathbf{V}$ for each attention head $h$. At each head $h$, $\mathbf{C}^{(h)}$ represents the dot product attention between queries $\mathbf{Q}^{(h)}$ and keys $\mathbf{K}^{(h)}$ scaled over dimension $d$ followed by softmax weighted aggregation over $\mathbf{V}^{(h)}$. $\mathbf{C}$ contains the aggregated outputs from all heads with final projection $\mathbf{W}^{O}$ applied. Using overlapping windows leads to spatially redundant time-aware tokens in $\mathbf{S}$, each with different local window context. Through Eq. 4, each redundant time-aware token in $\mathbf{S}$ is enabled to compare and weigh its contextually dissimilar variant, aligning them into a fused temporal representation. This allows $\mathbf{C}$ to yield rich time-aware clip level information by leveraging diverse temporal views from across overlapping windows. Recent works, e.g., [50], leverage clip sampler and clip encoder for long video understanding but their encoders are not timeaware and hence are not effective for temporal reasoning.
3.2 Bidirectional Long Short-Term Memory
Our time-sensitive clip encoder is applied on each short-term clip to obtain time-aware local features, which capture useful fine-grained cues for temporal reasoning within the short-term clip. However, they are not effective for temporal reasoning over the long-term video, which requires the ability to capture long-range temporal dependencies.
We introduce a BiLSTM module (i.e., green block in Fig. 2) for computing global features across multiple clips. We first concatenate time-aware local features extracted from clips in the temporal order which the clips appear in the video. We then pass the sequence of fused local features to a BiLSTM module for aggregating global features. Our BiLSTM module follows a standard architecture, including two LSTM networks: one processes the sequence in the original order (forward LSTM) and another processes the sequence in the reverse order (backward LSTM) as:

### v5

@@ Abstract
We introduce TemporalVLM, a video large language model (video LLM) for temporal reasoning and fine-grained understanding in long videos. Our approach includes a visual encoder for mapping a long-term video into features which are time-aware and contain both local and global cues. It first divides an input video into short-term clips, which are jointly encoded with timestamps and fused across overlapping temporal windows into time-sensitive local features. Next, the local features are passed through a bidirectional long short-term memory (BiLSTM) module for global feature aggregation. Moreover, to facilitate the evaluation of TemporalVLM, we present a large-scale long video dataset of industry assembly processes, namely IndustryASM, consisting of videos recorded on factory floors with actions and timestamps annotated by industrial engineers for time and motion studies and temporal action segmentation evaluation. Finally, extensive experiments show that TemporalVLM outperforms previous methods across temporal reasoning and fine-grained understanding tasks, i.e., dense video captioning, temporal video grounding, video highlight detection, and temporal action segmentation. To our best knowledge, our work is the first to incorporate LSTMs into video LLMs.
Video LLMs for Temporal Reasoning in Long Videos
Fawad J. Fateh    Umer Ahmed    Hamza Khan    M. Zeeshan Zia    Quoc-Huy Tran Retrocausal, Inc., Redmond, WA www.retrocausal.ai
@@ 1 Introduction
Figure 1: Video LLMs are often not time-sensitive (a, b), consider an input video as a single clip (a, c), and apply pooling (a, b) or query aggregation (c) for aggregating global features. Our model (d) includes a time-aware clip encoder for extracting time-aware fine-grained cues and a BiLSTM for capturing long-range temporal dependencies.
@@
Video temporal reasoning represents the process of reasoning about time and its passage in videos, with a focus on how events or actions happen and relate to each other in terms of time. Several video understanding applications require the ability of temporal reasoning, e.g., dense video captioning Wang et al. (2021); Zhu et al. (2022); Yang et al. (2023) and temporal action segmentation Li et al. (2020); Yi et al. (2021); Behrmann et al. (2022); Lu and Elhamifar (2024). Considerable efforts Yang et al. (2023); Luo et al. (2023a); Moon et al. (2023); Zhong et al. (2023); Lu and Elhamifar (2024); Lin et al. (2024) have been invested in developing models for solving individual tasks. These models often have different architectures. Thus, it is favorable to design a unified model for handling various tasks.
@@
The past few years have witnessed the impressive comprehension and generation capability of LLMs Achiam et al. (2023); Chiang et al. (2023); Taori et al. (2023); OpenAI (2024); Touvron et al. (2023a, b), which have emerged as a universal agent for performing various tasks. Video LLMs Li et al. (2023b); Liu et al. (2023); Luo et al. (2023b); Maaz et al. (2023); Zhang et al. (2023); Jin et al. (2024); Song et al. (2024) which incorporate video encoders with LLMs have been introduced. These methods often represent a video by a fixed number of tokens, yielding reduced performance with long videos, and encode frames and timestamps separately and hence struggle with temporal reasoning tasks. Please refer to the inferior results of these methods in Tab. 1. Recently, a few works Qian et al. (2024); Huang et al. (2024a, b); Ren et al. (2024) have developed video LLMs with temporal reasoning abilities, e.g., TimeChat Ren et al. (2024) proposes to vary the number of tokens based on the video length and jointly encode frames and timestamps. The above methods usually treat the entire video as a single clip Qian et al. (2024); Huang et al. (2024a, b); Ren et al. (2024) and aggregate tokens via pooling operation Huang et al. (2024b) and query aggregation Ren et al. (2024), struggling to capture fine-grained details in long videos.
@@
We propose TemporalVLM, a video LLM for temporal reasoning and fine-grained understanding in long videos. Fig. 1 shows the comparisons with prior works. Our model includes two architectural contributions. Firstly, we divide a long-term input video into multiple short-term clips and introduce a time-sensitive clip encoder for extracting fused time-aware local features from each clip. Secondly, we adopt a BiLSTM module which takes all the local features as inputs and computes global features from multiple clips. Our features not only are time-sensitive but also contain both local fine-grained and global semantic information, which are crucial for temporal reasoning in long videos. Moreover, to further evaluate our model, we present IndustryASM, a large-scale long video dataset of industry assembly processes for temporal action segmentation benchmarking and time and motion studies. Our IndustryASM dataset comprises of $4803$ videos with an average video duration of $119$ seconds. It covers in total $52$ diverse industry assembly tasks and includes timestamp and action labels. We convert the labels into chat samples with manually written instructions. Lastly, extensive experiments demonstrate that TemporalVLM achieves superior results over previous methods.
@@
  • We develop the first time-aware coarse-to-fine encoder, including a time-aware clip encoder (i.e., overlapping sliding video Q-Former) and a BiLSTM. By leveraging the modules, we tackle both fine-grained understanding and temporal reasoning in long videos.
  • We present IndustryASM, a large-scale long video dataset of manufacturing assembly procedures. Our videos are recorded on factory floors and labeled by industrial engineers.
  • TemporalVLM outperforms prior works on temporal reasoning and fine-grained understanding. To our best knowledge, this is the first work to blend LSTMs into video LLMs.
@@ 2 Related Work
Video Large Language Models. Video LLMs Li et al. (2023b); Liu et al. (2023); Luo et al. (2023b); Maaz et al. (2023); Zhang et al. (2023); Jin et al. (2024); Song et al. (2024) typically include a pre-trained visual encoder to extract visual features, a projection layer to map visual features into the text latent space of LLMs, and a pre-trained LLM for generating responses. They mostly differ in the visual encoder. VideoChat Li et al. (2023b) extracts frame features via a visual transformer Sharir et al. (2021) and employs a query transformer (Q-Former) Li et al. (2023a) to aggregate frame features into video features, while a video Q-Former is further included for temporal modeling in Video-LLaMA Zhang et al. (2023). The above methods usually map the video into a fixed number of tokens, yielding degrading performance with long videos, while encoding frames and timestamps separately and hence struggling with temporal reasoning. To address these drawbacks, methods with temporal reasoning capabilities are introduced, e.g., TimeChat Ren et al. (2024) presents a sliding video Q-Former to handle various video lengths and a time-aware frame encoder to jointly encode frames and timestamps. These methods often consider the video as a single clip Luo et al. (2023b); Maaz et al. (2023); Li et al. (2023b); Zhang et al. (2023) and aggregate tokens via pooling Luo et al. (2023b); Maaz et al. (2023) and query aggregation Li et al. (2023b); Zhang et al. (2023), overlooking fine-grained details. In this work, we divide the video into multiple clips and propose a time-aware clip encoder for capturing fused local features. Also, we integrate a BiLSTM module for aggregating global features.
@@
Video Temporal Reasoning. Temporal reasoning plays an important role in video understanding tasks, e.g., dense video captioning Wang et al. (2021); Zhu et al. (2022); Yang et al. (2023), temporal video grounding Wang et al. (2022); Luo et al. (2023a), video highlight detection Lei et al. (2021); Moon et al. (2023), and temporal action segmentation Li et al. (2020); Yi et al. (2021); Behrmann et al. (2022); Lu and Elhamifar (2024). Prior works Yang et al. (2023); Luo et al. (2023a); Moon et al. (2023); Lu and Elhamifar (2024) often focus on designing separate models for tackling individual tasks, while LLM-based models Qian et al. (2024); Huang et al. (2024a, b); Ren et al. (2024) capable of handling multiple tasks have emerged recently. Our TemporalVLM model belongs to the second group.
@@
Long Video Understanding. Challenges in long video understanding include complex spatial-temporal relationships and redundant information. Long video understanding methods have been developed via efficient architectures Donahue et al. (2015); Kondratyuk et al. (2021), temporal pooling/aggregation Sener et al. (2020); Wu and Krahenbuhl (2021), and clip selection Ghodrati et al. (2021); Gowda et al. (2021). For vision-language understanding tasks, methods based on temporal alignment Han et al. (2022); Buch et al. (2022) and memory Wu et al. (2022); Zhao et al. (2023) have been introduced. Recently, LongVLM Weng et al. (2024) divides the video into clips and employs a merging module for extracting local features and pooling operation for computing global features. It does not utilize timestamps, which are crucial for temporal reasoning. Our TemporalVLM model explicitly utilizes timestamps via a time-aware clip encoder. Moreover, we employ a learnable BiLSTM module for aggregating global features.
Procedural Activity Datasets. Existing datasets usually focus on cooking activities, e.g., Epic-Kitchens Damen et al. (2018), are curated from online sources and hence produced with multiple shots, e.g., COIN Tang et al. (2019), and are recorded with toy objects and lab environments, e.g., Assembly101 Sener et al. (2022). Our IndustryASM dataset focuses on manufacturing assembly processes, is recorded on factory floors, and is labeled by industrial engineers, thereby capturing procedural activities in realistic and practical environments that current datasets have not addressed.
3 TemporalVLM
Figure 2: TemporalVLM includes two novel components: a time-aware clip encoder for extracting time-aware fine-grained cues and a BiLSTM module for capturing long-range temporal dependencies.
An overview of TemporalVLM is shown in Fig. 2.
3.1 Short-Term Temporal Reasoning
@@
Video LLMs often consider an input video as a single clip and sample a fixed number of frames $N^{v}_{f}$ from the video, e.g., $N^{v}_{f}=96$ frames in TimeChat Ren et al. (2024). Moreover, prior works usually use query aggregation, e.g., TimeChat Ren et al. (2024), or pooling operation, e.g., Video-ChatGPT Maaz et al. (2023), to aggregate video tokens. Thus, they perform well with short videos but miss out fine-grained information for long videos.
@@
We introduce a time-aware clip encoder for extracting fused local fine-grained cues within a clip. We first divide the long-term input video into $C=6$ short-term clips and sample $N^{c}_{f}=96$ from each clip. Sampled frames of a clip along with timestamps are then passed to our time-aware clip encoder for jointly encoding frame contents and timestamps, yielding fused time-aware local features. Particularly, the time-aware frame encoder uses a pre-trained image encoder Sun et al. (2023) to obtain frame features, which are jointly encoded with timestamps via an image Q-Former Dai et al. (2023), yielding time-aware frame features $\mathbf{f}_{t}$.
Overlapping Sliding Video Q-Former. Frame features $\mathbf{f}_{t}$ are fed to a video Q-Former  Ren et al. (2024) across overlapping windows $\mathbf{W}_{i}$ of size $q$ = 32 frames and overlap $o$ = 16 frames. Window-wise outputs $\mathbf{V}_{i}$ are concatenated into $\mathbf{S}$ as:
$\mathbf{V}_{i}$ $=\text{Video Q-Former}(\mathbf{W}_{i}),$ (1)
$\mathbf{S}$ $=[\mathbf{V}_{1},\mathbf{V}_{2},\dots,\mathbf{V}_{W}],$ (2)
@@
with $W$ denoting the number of overlapping windows within a clip. We propose a fusion module to align diverse temporal cues from redundant boundary tokens in $\mathbf{S}$. Specifically, we apply multi-headed self-attention on $\mathbf{S}$, yielding $\mathbf{C}$ which fuses local contexts across multiple windows into a single context-aware embedding as:
@@
$\mathbf{C}$ $=[\mathbf{C}^{(1)},\mathbf{C}^{(2)},\dots,\mathbf{C}^{(H)}]\mathbf{W}^{O}.$ (4)
Here, we first project $\mathbf{S}$ into $\mathbf{Q}$, $\mathbf{K}$, $\mathbf{V}$ for each attention head $h$. At each head $h$, $\mathbf{C}^{(h)}$ represents the dot product attention between queries $\mathbf{Q}^{(h)}$ and keys $\mathbf{K}^{(h)}$ scaled over dimension $d$ followed by softmax weighted aggregation over $\mathbf{V}^{(h)}$. $\mathbf{C}$ contains aggregated outputs from all heads with final projection $\mathbf{W}^{O}$ applied. Using overlapping windows leads to spatially redundant time-aware tokens in $\mathbf{S}$, each with different local window contexts. This allows $\mathbf{C}$ to yield rich time-aware clip-level information by leveraging diverse temporal views from across overlapping windows. TimeChat’s Ren et al. (2024) video Q-Former uses non-overlapping windows and does not perform fusion, yielding inferior results as shown in Tab. 4.
3.2 Long-Term Temporal Reasoning
Our time-sensitive clip encoder is applied on each short-term clip to obtain fused time-aware local features, which capture useful fine-grained cues for temporal reasoning within the clip. However, they are not effective for temporal reasoning over the long-term video, which requires the ability to capture long-range temporal dependencies.
Bidirectional Long Short-Term Memory. We introduce a BiLSTM module for computing global features across multiple clips. We first concatenate time-aware local features extracted from clips in the temporal order the clips appear in the video. We then pass the sequence of local features to a BiLSTM for aggregating global features. Our BiLSTM follows a standard architecture, including two LSTM networks: one processes the sequence in the original order (forward) and another processes the sequence in the reverse order (backward) as: